A Bibliography of Publications in the Maple Newsletter and the Maple Technical Newsletter

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 September 2016 Version 1.09

Title word cross-reference

3 [Isb94]. 3n + 1 [Gon91]. 6 [Isb94]. D [MD92]. d_{mn}^{l}(β) [Isb94]. \sqrt{(}2) [GS93]. W [CGHJ93]. x^{2} ≡ a mod n [Vah93].

-j [Isb94].

2 [AL93, Mon88a].

3 [Mon94b]. 3D [HK97]. 3D-graphics [HK97].

4 [HT96, Lee96, Mon95a]. 4.0 [Ano87a, Mon88a]. 4.1 [Ano88e]. 4.2 [Ano88f]. 4.3 [Sco88].

5.3 [CTvB95].

Aberration [Dit96]. Aberrations [Mur97]. ACE [KLL^{+}97]. Acknowledgements [Ano92a, Ano92b, Ano93a, Ano93b, Ano94b]. Active [ER97]. Advances [Ano87a]. Aided [Mul96]. Air [NQC^{+}98]. Airborne [GN97]. Algae [AAK^{+}97]. Algebra [Ano88c, Hau94, Hau95, BH95, Man96, Mon93b]. Algebraic [KLL^{+}97, Ryb92, SKT96, FM91]. Algorithm [PVM97]. Algorithmic [MD92]. Algorithms [Pir95]. aliases [FM91]. American [Fee91, Gon88]. AMM [Fee89, Gon89a]. Analysis [AL93, BPS94, BPU98, Ged91, HRR96, Low98, MW96, Mea95, MD98, Par98, PL93, RS91, RR93, RF96]. Analytical [CTvB95, Mur97]. Analyze [GN97]. Analyzer [Rie96]. Analyzing [BC97, PVM97]. Anisotropic [MS98]. Announcement [Mon88a].

Announcements [Ano91, Ano92f, Ano92g, Ano93e, Ano93f, Ano94d, Fri89]. Answers [Ano87d, Ano88d]. Aplysia [ER97]. Application [Car94, Hof94, HB92, Par98, Pin94, PL93, Sco91a, STB97]. Applications [AL93, MD98, Mon94a, Pro97, SMFC91, SBG93, SFCM94]. Applied [Sto98, Gon88]. Approach [Sco88]. Approximate [Lop94a]. Approximation [Ged93]. Archimedes [Pat91b]. Areas [NQC^{+}98]. Arithmetic [BLM97]. Arithmetic-Geometric [BLM97]. Articles [Ano94a]. Assume [CM94]. Authors [Ano92c, Ano92d, Ano93c, Ano94a, Ano94c, Ano93d, Sco97a, Sco97b]. Automated [MW96]. Automatic [Gru97, Mon88b]. Availability [Ano87b, Ano88a, Ano88b]. AVS [Ano92e].

Band [Sac97]. based [AS98]. Bases [Gru93b, RF96]. Basic [Mul96]. Behavior [EW97]. Bernoulli [Mea95]. Bessel [MDS92]. Bifurcation [CC97]. Billiard [GG92]. Binary [Tay94]. Binomial [Str93]. Biochemistry [GJS94]. Biology [AAK^{+}97, HY94]. Bisection [Gru97]. Boundary [MHW96]. Brief [Ano87c]. Broad [SBG93]. Broken [Pat91b]. Building [Gon92].

C [CG96]. Calculation [SRGH95]. Calculator [BL92]. Calculus [BDL95]. Cam [AS98, PF96]. Capabilities [Ryb92]. case [GM88]. Center [Pro97]. Chaos [Gir97]. Characteristics [Rus94]. Chemical [HB92, Par98, SH94]. Chord [Pat91b]. Circles [PWMdG91, WMP94]. Circuit [Rie96]. Class [AAK^{+}97]. Classes [SBG93]. Cloth [EW97]. Code [Ayo95, CG96, Len98, SGMS97]. Codes [GVC95]. Coefficients [Ham95]. Collect [Mon89a]. Combinatorial [Par98, Zim94]. Combinatorics [Fei94, KLL^{+}97]. Combining [Bac97]. Competition [AAK^{+}97, DS95]. Computation [Mon92, OM97, Ogu96]. Computations [Gon91]. Computer [Ano88c, Mon93b, Mul96]. Computing [GS93, RSPL98, Ryb92]. Conceptual [RSPL98]. Conductor [SRGH95]. Congruence [Vah93]. Conjecture [Gon91, Lis94]. Construction [KS93]. Contagion [GN97]. Control [Ogu96, PL93]. Controlled [RR93]. Controlling [Sto98]. Convinced [Fri89]. Cooke [Dit96]. Counting [Gui92]. Course [BH95]. Cray [Mon88a]. Curves [Mul96].

Decay [Far95]. defined [FM91]. Definite [Sco88, Sco91b, SBG93]. Design [AS98, AL93, Mul96, PF96, PL93, RSPL98]. Development [Pin94]. Differential [BM97b, CTvB95, Cyg96, Fee91, GM97, Her97, Lab95, Len98, Lop94b, AL95, Man96, SKT96]. Differentiation [Gru97, Jég97, MD92]. diffgrob2 [Man96]. Diffusion [Ham95]. digits [GS93]. dimensional [Gir97]. Diophantine [Ogu98]. Discrete [Ogu98]. Discrete-Time [Ogu98]. Distribution [SRGH95]. Distributions [AL95]. Disturbing [HM94]. Draping [EW97]. Drives [BPS94]. dynamical [Gir97]. Dynamics [MW96, Sch92].

E2983 [Fee89]. E3229 [Gon89a].

Econometrics [Gre94]. Economy [CC97]. Editors [BM97a, HM97, OC98, SR92a, SR92b, SR93, SP93, SM94, SP94]. Education [Ano88c, SPR92, SMPR94, SMC94]. Eigensystems [BH95]. Eigenvalue [Mon92]. Eigenvalues [Mon93b]. Eigenvector [Mon92]. Eigenvectors [Mon93b]. Einstein [Cza95]. Engine [Mon95a]. Engineering [AS98, AL93, Cor96, LK96, RM96, SMFC91, SFCM94]. Engineers [EM97, Lee96]. Epicyclic [BPS94]. Equation [Fee91, Ham95, Lop94b, Sac97]. Equations [AT94, BM97b, Bri95, CTvB95, Cyg96, Cza95, Dit96, GJS94, GM97, Her97, Lab95, Len98, AL95, Mea95, Ogu98, SH94, TM93, SKT96, Aru97]. Equilibria [Tay94]. Estimation [BK96, AL95]. Evaluation [Far95, Gra92]. Evaporator [KTT98]. Exact [Cza95, Lop94a, Mon92]. Examples [Gon88]. Expansions [HM94]. Exploratory [MD98]. Exponentials [Sco91b]. Exponents [GM88]. Extended [Ryb92].

Facility [CM94, Mon88d]. Factor [Far95]. factorization [FM91]. Factorizing [Mel97]. Features [Ano91, Lee96]. Fermat [GM88, GM94]. Field [Moh95, STB97]. fields [FM91]. finite [FM91]. First [GM88]. Flexible [AS98, Pie96]. Flow [BPS94]. Flowsheet [Par98]. Form [Far95]. Formal [Bri95, GK95]. Formula [Gru93a, Sal91]. FORTRAN [SGMS97, CG96, Mon89b]. Foxes [Sha97]. Frenet [Aru97]. Frenet-Serret [Aru97]. Frequency [Low98]. Full [Dev89]. Function [CGHJ93, HM94, MDS92, Mon89a]. Functional [Bri95]. Functions [Pir95, Ryb92, SGMS97, Isb94, Sac97].

Gaïa [Zim94]. Gauss [Gru93b, GM93]. Gaussian [SGMS97]. Gaussian-Type [SGMS97]. Gear [BPS94]. General [Car94]. Generating [Hau94, Hau95]. Generation [Ayo95, CG96, KG94, Len98, SGMS97, Zim94]. Genetics [AL95]. Geometric [BLM97, Mul96]. Geometry [Sto98]. get [Sal91]. Glasses [Dit94]. Graphical [HT96, Mea95]. graphics [HK97, LM94]. Graphs [Gui92]. Gröbner [Gru93b, RF96]. Group [Wal97]. Groups [Fei94]. Guidelines [AS98, Ano92c, Ano92d, Ano93c, Ano94c, Ano93d, Sco97a, Sco97b].

Harmonic [RR93]. Hashing [PVM97].

Heat [Ham95]. Heavy [Far95]. Help [Mon93b, PVM97]. Heuristic [Gon89b]. High [BL92]. History [Ano87c, WMP94]. HIV [Her93]. Hopf [CC97]. Hypersingular [Gra92].

Identification [Ham95]. Identities [Str93]. II [Hau95, Sco91b, SMPR94, Tay94]. III [AT94]. Impact [Mea97]. Implementation [KTT98, BDL95, LM94]. Indefinite [Pat91a, Pir95]. Industrial [KTT98]. Infection [Her93]. Infinite [Gru94, Mel97]. Insights [Lop94b]. Instruction [LK96]. Instructors [Lop96]. Integral [Dev88a, Gra92, MDS92]. Integrals [Sco91b, SBG93, SGMS97, HK97]. Integrating [LK96]. Integration [Ged91, Sco88]. Interface [HT96, SH94]. Intersecting [Sto98]. Intervals [GVC95]. Introduction [GM93]. Introductory [GM97, BH95]. Inverse [Jég97]. Inversion [MS98]. Investigating [BH95]. Involving [Sco91b]. Isolating [GVC95]. Isomers [Fei94]. Isomorphic [Gui92]. iterated [HK97]. IV [Tay96].

j [Isb94].

Kinematic [BPS94, BPU98, RF96].

Kinetic [SH94]. Kinetics [HB92]. Kostant [Lis94].

Lambert [CGHJ93]. Lamé [Sac97]. Last [GM88, GM94]. Layer [SRGH95]. LCFS [PVM97]. Lens [Dit96]. Library [Mon94a]. Likelihood [BK96]. linalg [Joh92]. Linear [Ayo95, Hau94, Hau95, BH95, Lab95, PVM97, AL95]. Logarithms [Sco91b].

Mac [Dev89]. Machines [AL93]. MacroC [CG96]. Macrofort [CG96]. Manifolds [Pro97]. Manipulating [Mon88c]. Manipulation [GVC95, Gru93a]. Maple [HM97, PVM97, Cyg96, KLL^{+}97, AT94, AS98, Ano87a, Ano87b, Ano87c, Ano88a, Ano88b, Ano88e, Ano88f, Ano91, Ano92e, AL93, Ayo95, BC97, BL92, BM97b, BPU98, CTvB95, CG96, Car94, CGHJ93, Cor96, Cza95, Dev88a, Dev89, Dit94, ER97, EM97, GN97, Gon88, Gra92, Gru93a, Gru94, Gui92, GM97, Ham95, HY94, Her97, Hof94, HB92, HT96, Isb94, Jég97, KS93, GK95, Koe96, Kra94, Lab95, Lee96, Lis94, Lop96, LM94, Mel97, MD98, Moh95, Mon88a, Mon88b, Mon88c, Mon88d, Mon89a, Mon92, Mon93a, Mon94a, Mon94b, Mon94c, Mon94d, Mon95a, Mon95b, Mon95c, Mur97, Ogu96, Ogu98, Pat91a, SPR92, PF96, Pie96, Pin96, Pin94, Pir95, PL93, RS92, RM96, Sal91, DS95, Sco88, SMFC91, Sco91a, SFCM94, SMPR94, SMC94, STB97, Sha97, SH94]. Maple [TM93, Tay94, SKT96, Tay96, KTT98]. Maple-based [AS98]. Math [Mon95a]. MathEdge [LK96, Pin94]. Mathematical [AAK^{+}97, Fee91, Gon88, SMFC91, SFCM94]. Mathematics [Cor96, HY94]. Matrix [Ogu98, Isb94]. Maximum [BK96]. Maxwell [AT94]. McConnell [GJS94]. Mean [Dev88b]. Means [BLM97, BPU98]. Mechanical [AS98, MW96, Sco91a]. Mechanics [Isb94]. Mechanisms [BPU98, PF96, RF96]. Meson [Far95]. Method [BC97, Low98]. mid [GM94]. million [GS93]. Model [CC97, GN97, Her93]. Modeling [EW97, ER97, Hof94, Kra94, Mea97, Moh95]. Modelling [Pie96]. Models [Ayo95]. Molecular [OM97, SGMS97]. Mollusc [ER97]. Monthly [Fee91, Gon88]. Mple [Mea97]. Multi [KTT98]. Multi-stage [KTT98]. Multimedia [LK96]. Multiprocessing [RS91]. Multirate [HRR96].

NAG [ER97]. National [CC97]. Nature [Mon93b]. Near [STB97]. Near-Field [STB97]. Network [Bac97]. Neuron [ER97]. News [Ano91, Ano92f, Ano92g, Ano93e, Ano93f, Ano94d, Fri89]. No [BM97a, HM97]. Non [Ayo95, Gui92, AL95]. Non-Isomorphic [Gui92]. Non-linear [Ayo95, AL95]. Nonlinear [EM97]. Note [BM97a, HM97, SR92a, SR92b, SR93, SP93, SM94, SP94, Sha97, OC98]. Number [KG94, FM91]. Numbers [Ryb92]. Numeric [Bac97]. Numerical [Ayo95, Fee91, Ged91, Ged93, KS93, Mea95, RSPL98]. Numerically [SH94].

Objectives [Dit94]. One [Gir97, GS93]. One-dimensional [Gir97]. Operator [MD92]. Optical [Mur97]. Optics [STB97]. Optimal [PL93]. Optimization [Bac97, Mon89b]. Optimized [Low98, SGMS97]. Oscillator [Lop94a]. Overview [Ogu96].

Package [CTvB95, Ged93, GK95, BDL95, Man96, Zim94]. Packages [Dev88a]. Packing [PWMdG91, WMP94]. Pairfunctions [Wal97]. Parachute [Mea97]. Parameter [MS98]. Parameters [BK96]. Part [Dev88a, Sco91b, SMPR94]. Partial [CTvB95]. Phase [RR93, Sch92, Tay94]. Phylogenetic [Gon92]. Physical [HM97]. Physics [EM97]. Planning [MD98]. Plants [Par98]. Plot [BK96]. Plotting [Mon88d]. Plus [Dev89]. point [MHW96]. Pollution [NQC^{+}98]. Polynomials [GVC95, Mon88c]. Population [AL95, Sch92]. Portraits [Sch92]. Potential [SRGH95]. Power [Bac97, BPS94, Bri95, Dev89, Dit96, GK95]. Power-Flow [BPS94]. Precision [BL92]. presented [Fee91]. Primality [Gon89b]. Prime [GM88]. Principles [Sco91a]. Probability [BK96]. Probing [PVM97]. Problem [Fee89, GG92, Gon89a, Gon92, Sco91a, Fee91, Mea97]. Problems [Hau94, Hau95, Jég97, MHW96, SMFC91, SFCM94, Gon88]. Procedures [KTT98]. Process [Ogu96, RSPL98]. Programmers [Mon95b, Mon95c]. Project [AAK^{+}97].

Projection [BC97]. Proof [Dev88b]. Propagation [NQC^{+}98]. Properties [Tay96]. Property [AT94]. Putnam [DS95].

Quantum [Isb94, Sco91a]. Questions [Ano87d, Ano88d]. Queuing [RS91].

R15 [ER97]. Random [KG94, Zim94]. Rational [Pir95]. Rayleigh [Low98]. Real [GVC95]. Recipes [Sco91b, SBG93]. Relations [AT94, Wal97]. Relativity [Car94]. Release [AL93, HT96, Lee96, Mon94b, Mon95a]. Resource [AAK^{+}97]. Revisited [BLM97]. Revolution [RS92]. Rhythmically [ER97]. Robotic [Sto98]. Robots [Kra94, Pie96]. Roots [GVC95]. rotation [Isb94].

Saw [Lop94a]. Saw-Toothed [Lop94a]. Schemes [KS93]. Science [SMC94]. Sciences [HM97]. Scientific [SMFC91, SFCM94]. Scientists [EM97, Lee96]. section [Fee91]. Seismology [MS98]. Selection [Dit94]. Semigroup [Wal97]. Sensitivity [MS98]. Separable [Lop94b]. Series [Bri95, GK95]. Serret [Aru97]. Share [Mon94a]. Shooting [MHW96]. Short [BM97b, Sha97]. Shotgun [Rus94]. Simple [Cor96, Sco91a]. Simplification [CM94, Mon88b]. Simulating [KTT98, Kra94]. simulations [Gir97]. Snow [Hof94]. Solution [Bri95, Fee91, Len98, MHW96, Ogu98]. Solutions [Cza95, Lop94a]. Solve [Jég97]. Solving [CTvB95, Cyg96, Dit96, GJS94, Her97, Lab95, Lis94, Sco91b, SBG93, SH94, SKT96, Vah93]. Source [Ayo95]. Spatial [BPU98]. Spectroscopy [OM97]. Spherical [SRGH95]. Square [PWMdG91, WMP94]. stage [KTT98]. State [TM93]. Status [GM94]. Steam [Tay96]. Stiff [SKT96]. Stirling [Sal91]. Stochastic [Cyg96, KS93]. Structures [Gru94, Zim94]. Students [RM96]. Study [Mur97]. Summation [Koe96, Pat91a, Pir95]. Sums [Str93]. Surfaces [Mul96, RS92]. Symbolic [Bac97, Dev88b, Ged91, HRR96, Jég97, Low98, MW96, Mea95, OM97, Ogu96, RSPL98, Rie96]. symbols [Isb94]. Symmetric [Ogu98]. Synthesis [Par98]. Syrup [Rie96]. System [Bac97, MW96, Sco91a]. Systems [Ano88c, HRR96, AL95, Low98, PL93, Tay94, SKT96, Gir97].

Teaching [GM97, RM96]. Technique [MHW96]. Techniques [RSPL98]. Telescope [Dit94]. Testing [Gon89b, KG94]. Their [Ano94a]. Theorem [Dev88b, GM94, Pat91b, GM88]. Thermodynamic [AT94]. Thermodynamics [AT94, TM93, Tay94, Tay96]. Thom [GVC95]. Three [SRGH95, Sal91]. Three-Layer [SRGH95]. Time [Ogu98]. Tips [Lop96, Mon93a, Mon94c, Mon94d, Mon95b, Mon95c]. Tool [AS98]. Toolkit [Pin94]. Tools [Bac97]. Toothed [Lop94a]. Torch [Sto98]. transition [Gir97]. Translation [Mon89b]. Transport [Hof94]. Transportation [MD98]. Traveltime [MS98]. Tree [Gon92]. Trees [Sha97]. Triplet [Dit96]. Tubes [Sto98]. Turtle [LM94]. Tutorial [BM97b, Her97]. Two [MHW96]. Two-point [MHW96]. Type [SGMS97].

Umbral [BDL95]. Unary [Wal97]. Understand [Mon93b, Sch92]. Urban [NQC^{+}98]. User [FM91, HT96]. Users [Mon93a, Mon94c, Mon94d, Mon95b, Mon95c]. Using [Dit94, GN97, Her97, Jég97, Lis94, MD98, Mon93b, Mur97, PF96, RF96, Sch92, Ayo95, ER97, Ged91, Gra92, Gru93a, Ham95, Mel97, Moh95, Ogu98, RSPL98, Sco91a].

V [AL93, HT96, Lee96, Mon94b, Mon95a, Ano91, Ano92e, LM94, Ogu96, Ogu98, Tay94, SKT96]. Value [Dev88b, MHW96]. Variational [Sco91a]. Verification [Wal97]. Vibratory [Low98]. Visualizing [Aru97]. Vol [BM97a, HM97]. Volume [SRGH95]. Volumes [RS92].

Waveforms [RR93]. Wavefront [Mur97]. Wavelets [Gre94]. ways [Sal91]. Web [Pin96]. Weibull [BK96]. Welding [Sto98]. Wide [Pin96]. Winter [BM97a]. within [CG96]. words [Mel97]. Working [Joh92]. World [Pin96]. Wounding [Rus94]. Woven [EW97].

References

Agler:1997:RCA

[AAK^{+}97] Heather Agler, Aaron Ahearn, Anne Kitchell, Nicole Lopanik, Heather Miller, Matthew Miller, and David S. Wethey. Resource competition in algae: A class project in mathematical biology. Maple Technical Newsletter, 4 (2):78–85, 1997. ISSN 1061-5733.

Argent:1993:EAM

[AL93] Jason Argent and Tom Lee. Engineering applications of Maple V Release 2: Analysis and design of machines. Maple Technical Newsletter, 0(9):54– 61, Spring 1993. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn9.html.

Lopez:1995:PGE

[AL95] N. E. Abrouk and R. J. Lopez.

Population genetics: Estimation of distributions through systems of non-linear differential equations. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n1.html.

Anonymous:1987:AMa

[Ano87a] Anonymous. Advances with Maple 4.0. Maple Newsletter, 0(1):??, January 1987. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn1.html.

Anonymous:1987:AMb

[Ano87b]  Anonymous.  Availability

of Maple. Maple Newsletter, 0(1):??, January 1987. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn1.html.

Anonymous:1987:BHM

[Ano87c] Anonymous. A brief history of Maple. Maple Newsletter, 0(1):??, January 1987. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn1.html.

Anonymous:1987:QA

[Ano87d] Anonymous. Questions and answers. Maple Newsletter, 0(1):??, January 1987. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_

Packages/Per_Purpose/General/
Maple/mtn/mtn1.html.

Anonymous:1988:AMa

[Ano88a]  Anonymous.  Availability

of Maple. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn2.html.

Anonymous:1988:AMb

[Ano88b] Anonymous. Availability of Maple. Maple Newsletter, 0 (3):??, July 1988. ISSN 1074-

3790. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn3.html.

Anonymous:1988:CAS

[Ano88c] Anonymous. Computer algebra systems in education. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-3790. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn2.html.

Anonymous:1988:QA

[Ano88d] Anonymous. Questions and answers. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn2.html.

Anonymous:1988:WNMa

[Ano88e] Anonymous. What’s new in Maple 4.1. Maple Newsletter, 0(2):??, January 1988.

ISSN 1074-3790. URL http:

                                        //www.can.nl/Systems_and_
                                        Packages/Per_Purpose/General/
                                        Maple/mtn/mtn2.html.

Anonymous:1988:WNMb

[Ano88f] Anonymous. What’s new in Maple 4.2. Maple Newsletter, 0 (3):??, July 1988. ISSN 1074-

                                        3790. URL http://www.can.
                                        nl/Systems_and_Packages/Per_
                                        Purpose/General/Maple/mtn/
                                        mtn3.html.

Anonymous:1991:NAF

[Ano91] Anonymous. News and announcements: Features of Maple V. Maple Technical Newsletter, 0(5):??, Winter 1991. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn5.html.

Anonymous:1992:Aa

[Ano92a]  Anonymous.  Acknowledge-

ments. Maple Technical Newsletter, 0(7):56, Spring 1992. ISSN 1061-5733.

Anonymous:1992:Ab

[Ano92b]  Anonymous.  Acknowledge-

ments. Maple Technical Newsletter, 0(8):61, Fall 1992. ISSN 1061-5733.

Anonymous:1992:GAa

[Ano92c] Anonymous. Guidelines for authors. Maple Technical Newsletter, 0(7):56, Spring 1992. ISSN 1061-5733.

Anonymous:1992:GAb

[Ano92d] Anonymous. Guidelines for authors. Maple Technical Newsletter, 0(8):61, Fall 1992. ISSN 1061-5733.

Anonymous:1992:MVA

[Ano92e] Anonymous. Maple V and AVS.

Maple Technical Newsletter, 0 (7):9–10, Spring 1992. ISSN 1061-5733.

Anonymous:1992:NAa

[Ano92f] Anonymous. News and announcements. Maple Technical Newsletter, 0(7):4–8, Spring 1992. ISSN 1061-5733.

Anonymous:1992:NAb

[Ano92g] Anonymous. News and announcements. Maple Technical Newsletter, 0(8):4–15, Fall 1992. ISSN 1061-5733.

Anonymous:1993:Aa

[Ano93a]  Anonymous.  Acknowledge-

ments. Maple Technical Newsletter, 0(9):78, Spring 1993. ISSN 1061-5733.

Anonymous:1993:Ab

[Ano93b]  Anonymous.  Acknowledge-

ments. Maple Technical Newsletter, 0(10):94, Fall 1993. ISSN 1061-5733.

Anonymous:1993:GA

[Ano93c] Anonymous. Guidelines for authors. Maple Technical Newsletter, 0(9):77, Spring 1993. ISSN 1061-5733.

Gruntz:1993:FMUb

[Ano93d] Anonymous. Guidelines for authors. Maple Technical Newsletter, 0(10):93, Fall 1993. ISSN 1061-5733.

Anonymous:1993:NAa

[Ano93e] Anonymous. News and announcements. Maple Technical Newsletter, 0(9):4–11, Spring 1993. ISSN 1061-5733.

Anonymous:1993:NAb

[Ano93f] Anonymous. News and announcements. Maple Technical Newsletter, 0(10):6–9, Fall 1993. ISSN 1061-5733.

Anonymous:1994:AAT

[Ano94a] Anonymous. About the articles and their authors. Maple Technical Newsletter, 0:2–5, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Anonymous:1994:A

[Ano94b]  Anonymous.  Acknowledge-

ments. Maple Technical Newsletter, 1(1):94, Spring 1994. ISSN 1061-5733.

Anonymous:1994:GA

[Ano94c] Anonymous. Guidelines for authors. Maple Technical Newsletter, 1(1):93, Spring 1994. ISSN 1061-5733.

Anonymous:1994:NA

[Ano94d] Anonymous. News and announcements. Maple Technical Newsletter, 1(1):4–6, Spring 1994. ISSN 1061-5733.

Aruliah:1997:VFS

[Aru97] D. A. Aruliah. Visualizing the Frenet-Serret equations. Maple Technical Newsletter, 4(1):105– 110, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Alpers:1998:FME

[AS98] Burkhard Alpers and Stephan Steeb. Flexible mechanical engineering guidelines — A Maplebased tool for cam design. Maple Technical Newsletter, 5(2–3):40– 48, 1998. ISSN 1061-5733.

Adams:1994:TMI

[AT94] Steve A. Adams and Ross Taylor. Thermodynamics with Maple: III thermodynamic property relations and the Maxwell equations. Maple Technical Newsletter, 1(2):68–81, Fall 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n2.

html. See also [TM93, Tay94, Tay96].

Ayotte:1995:SCG

[Ayo95] K. W. Ayotte. Source code generation for linear and nonlinear numerical models using Maple. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n1.html.

Bacher:1997:CSN

[Bac97] Rainer Bacher. Combining symbolic and numeric tools for power system network optimization. Maple Technical Newsletter, 4(2):41–51, 1997. ISSN 1061-5733.

Bauschke:1997:APM

[BC97] Heinz H. Bauschke and Robert M. Corless. Analyzing a projection method with Maple. Maple Technical Newsletter, 4(1):2–7, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Loeb:1995:IUC

[BDL95] A. Bottreau, A. Di Bucchianico, and D. Loeb. Implementation of an umbral calculus package. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v2n2.html.

Hirst:1995:IEI

[BH95] W. C. Bauldry and H. P. Hirst. Investigating eigensystems in an introductory linear algebra course. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v2n2.html.

Bohoris:1996:WPP

[BK96] G. Bohoris and P. A. Kostagiolas. Weibull probability plot and maximum likelihood estimation of its parameters. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Borwein:1992:MHP

[BL92] Jonathan M. Borwein and Mark A. Limber. Maple as a high precision calculator. Maple Technical Newsletter, 0(8):39–44, Fall 1992. ISSN 1061-5733. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn8.html.

Borwein:1997:AGM

[BLM97] Jonathan A. Borwein, Petr Lisonek, and John A. Macdonald. Arithmetic-geometric means revisited. Maple Technical Newsletter, 4(1):20–27, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Borwein:1997:NEW

[BM97a] Jon Borwein and Michael Monagan. A note from the Editors: Winter 1997, vol. 4, no.

1. Maple Technical Newsletter, 4(1):1, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Borwein:1997:DEM

[BM97b] Jonathan Borwein and Michael Monagan. Differential equations in Maple: A short tutorial. Maple Technical Newsletter, 4(1):55–58, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Belfiore:1994:KPF

[BPS94] Nicola Pio Belfiore, Ettore Pennestri, and Rosario Sinatra. Kinematic and powerflow analysis of epicyclic gear drives. Maple Technical Newsletter, 1(1):47–55, Spring 1994. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtnv1n1.
                                        html.

Brutti:1998:KAS

[BPU98] Carlo Brutti, Ettori Pennestrì, and Fabrizio Urbinati. Kinematic analysis of spatial mechanisms by means of Maple. Maple Technical Newsletter, 5(2–3):49– 57, 1998. ISSN 1061-5733.

Briggs:1995:FPS

[Bri95]  K. Briggs.  Formal power

series solution of functional equations. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v2n2.html.

Carminati:1994:AMG

[Car94] J. Carminati. Application of Maple to general relativity. Maple Technical Newsletter, 0: 43–48, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Calpaert:1997:HBM

[CC97] Jan Calpaert and Douglas Curran. A Hopf bifurcation in a model of a national economy. Maple Technical Newsletter, 4 (1):88–91, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Capolsini:1996:MMC

[CG96] P. Capolsini and C. Gomez.

MacroC and Macrofort: C and Fortran code generation within Maple. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Corless:1993:LFM

[CGHJ93] Robert M. Corless, Gaston H.

Gonnet, D. E. G. Hare, and David J. Jeffrey. Lambert’s W function in Maple. Maple Technical Newsletter, 0(9):12– 22, Spring 1993. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn9.html.

Corless:1994:SAF

[CM94] Robert M. Corless and Michael B.

Monagan. Simplification and the Assume facility. Maple Technical Newsletter, 1(1):24– 31, Spring 1994. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_

                                        Purpose/General/Maple/mtn/
                                        mtnv1n1.html.

Corless:1996:SEM

[Cor96] R. M. Corless. Simple engineering mathematics with Maple. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Bulow:1995:MPA

[CTvB95] E. S. Cheb-Terrab and K. von Bülow. A Maple 5.3 package for the analytical solving of partial differential equations. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733.

                                        URL http://web.mit.edu/
                                        afs/athena.mit.edu/software/
                                        maple/www/plibrary/mtn/mtn-
                                        v2n2.html.

Cyganowski:1996:SSD

[Cyg96] S. Cyganowski. Solving stochastic differential equations with MAPLE. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733.

                                        URL http://web.mit.edu/
                                        afs/athena.mit.edu/software/
                                        maple/www/plibrary/mtn/mtn-
                                        v3n2.html.

Czapor:1995:MES

[Cza95] S. R. Czapor. Maple and exact solutions of Einstein’s equations. Maple Technical Newsletter, 2(2):??, Fall 1995.

ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n2.html.

Devitt:1988:PIP

[Dev88a] J. S. Devitt. Packages as an integral part of Maple. Maple Newsletter, 0(3):??, July 1988. ISSN 1074-3790. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn3.html.

Devitt:1988:SPM

[Dev88b] J. S. Devitt. A symbolic proof of the mean value theorem. Maple Newsletter, 0(3):??, July 1988. ISSN 1074-3790. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn3.html.

Devitt:1989:FPM

[Dev89] J. S. Devitt. The full power of Maple on a Mac Plus. Maple Newsletter, 0(4):??, January 1989. ISSN 1074-3790. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn4.html.

Ditteon:1994:UMS

[Dit94] Richard Ditteon. Using Maple in the selection of glasses for telescope objectives. Maple Technical Newsletter, 1(2):62–67, Fall 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n2.
html.

Ditteon:1996:SAP

[Dit96]  R. Ditteon.  Solving the

aberration and power equations for the Cooke triplet lens. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v3n2.html.

Salvy:1995:MPC

[DS95] P. Dumas and B. Salvy. Maple and the Putnam Competition. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v2n2.html.

Enns:1997:NPM

[EM97] Richard H. Enns and George C. McGuire. Nonlinear physics with Maple for scientists and engineers. Maple Technical Newsletter, 4(2):93–99, 1997. ISSN 1061-5733.

Elliott:1997:MRA

[ER97] C. J. H. Elliott and J. S. Robinson. Modeling the rhythmically active neuron R15 of the mollusc Aplysia using Maple and NAG. Maple Technical Newsletter, 4(2):6–14, 1997. ISSN 1061-5733.

Eberhart:1997:MDB

[EW97] Bernd Eberhart and Andreas Weber. Modeling the draping behavior of woven cloth. Maple Technical Newsletter, 4(2):25– 31, 1997. ISSN 1061-5733.

Fariborz:1995:FFE

[Far95] A. H. Fariborz. A form factor evaluation in heavy meson decay. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n2.html.

Fee:1989:APE

[Fee89]  G. J. Fee.  AMM prob-

lem E2983. Maple Newsletter, 0(4):??, January 1989. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn4.html.

Fee:1991:NSD

[Fee91] Greg Fee. Numerical solution of a differential equation presented in the problem section of the American Mathematical Monthly. Maple Technical Newsletter, 0(5):??, Winter 1991. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn5.html.

Feigelstock:1994:IGC

[Fei94]  Shalom Feigelstock.  Iso-

mers, groups, and combinatorics. Maple Technical Newsletter, 1(2):82–85, Fall 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/

General/Maple/mtn/mtnv1n2.
html.

Fee:1991:UDA

[FM91] Gregory Fee and Michael Monagan. User defined aliases and factorization over finite and algebraic number fields. Maple Technical Newsletter, 0(5):??, Winter 1991. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn5.html.

Friel:1989:NAN

[Fri89] J. William Friel. News and announcements: Now, are you convinced? Maple Newsletter, 0(4):??, January 1989. ISSN 1074-3790. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn4.html.

Geddes:1991:NIU

[Ged91]  Keith O. Geddes.  Numer-

ical integration using symbolic analysis. Maple Technical Newsletter, 0(6):??, Fall 1991. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn6.html.

Geddes:1993:PNA

[Ged93]  K. O. Geddes.  A pack-

age for numerical approximation. Maple Technical Newsletter, 0(10):28–36, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Gander:1992:BP

[GG92] Walter Gander and Dominik Gruntz. The billiard problem. Maple Technical Newsletter, 0(7):44–49, Spring 1992. ISSN 1061-5733. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn7.html.

Giraldo:1997:ODD

[Gir97] Antonio Giraldo. One-dimensional dynamical systems. dynamical simulations of the transition to Chaos. Maple Technical Newsletter, 4(1):99–104, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Grotendorst:1994:SME

[GJS94] Johannes Grotendorst, Paul Jansen, and Siegfried M. Schoberth. On solving the McConnell equations in biochemistry. Maple Technical Newsletter, 1(1):63–68, Spring 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n1.
html.

Koepf:1995:MPF

[GK95] D. Gruntz and W. Koepf.

Maple package on formal power series. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n2.html.

Granville:1988:FCF

[GM88] Andrew Granville and Michael Monagan. The first case of Fermat’s Last Theorem for all prime exponents up to 714,591,416,091,389. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-3790. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn2.html.

Gruntz:1993:IG

[GM93] D. Gruntz and M. Monagan. Introduction to Gauss. Maple Technical Newsletter, 0(9):23– 35, Spring 1993. ISSN 1061-

                                        5733. URL http://www.can.
                                        nl/Systems_and_Packages/Per_
                                        Purpose/General/Maple/mtn/
                                        mtn9.html.

Granville:1994:SFL

[GM94] Andrew Granville and Michael Monagan. The status of Fermat’s last theorem — mid 1994. Maple Technical Newsletter, 0: 6–9, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Guidera:1997:TID

[GM97] Carollyne Guidera and Michael Monagan. Teaching introductory differential equations with Maple. Maple Technical Newsletter, 4(1):59–67, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Gammaitoni:1997:UMA

[GN97] L. Gammaitoni and M. C.

Nucci. Using Maple to analyze a model for airborne contagion. Maple Technical Newsletter, 4 (2):2–5, 1997. ISSN 1061-5733.

Gonnet:1988:EMA

[Gon88]  G. H. Gonnet.  Examples

of Maple, applied to problems from the American Mathematical Monthly. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-3790. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn2.html.

Gonnet:1989:APE

[Gon89a] Gaston H. Gonnet. AMM problem E3229. Maple Newsletter, 0(4):??, January 1989. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn4.html.

Gonnet:1989:HPT

[Gon89b] Gaston H. Gonnet. Heuristic primality testing. Maple Newsletter, 0(4):??, January 1989. ISSN 1074-3790. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn4.html.

Gonnet:1991:CC

[Gon91]  Gaston H. Gonnet.  Com-

putations on the 3n + 1 conjecture. Maple Technical Newsletter, 0(6):??, Fall 1991. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn6.html.

Gonnet:1992:PPT

[Gon92] Gaston H. Gonnet. A problem in phylogenetic tree building. Maple Technical Newsletter, 0(7):31–38, Spring 1992. ISSN 1061-5733. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn7.html.

Gray:1992:HIE

[Gra92]  L. J. Gray.  Hypersingu-

lar integral evaluation using Maple. Maple Technical Newsletter, 0(8):57–60, Fall 1992. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn8.html.

Greenblatt:1994:WE

[Gre94] Seth A. Greenblatt. Wavelets in econometrics. Maple Technical Newsletter, 0:10–16, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Gruntz:1993:FMUa

[Gru93a] Dominik Gruntz. Formula manipulation using Maple. Maple Technical Newsletter, 0(10):85– 92, Fall 1993. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn10.html.

Gruntz:1993:GBG

[Gru93b] Dominik Gruntz. Gröbner bases in Gauss. Maple Technical Newsletter, 0(9):36–46, Spring 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn9.html.

Gruntz:1994:ISM

[Gru94] Dominik Gruntz. Infinite structures in Maple. Maple Technical Newsletter, 1(2):19–30, Fall 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n2.
html.

Gruntz:1997:ADB

[Gru97] Dominik Gruntz. Automatic differentiation and bisection. Maple Technical Newsletter, 4 (1):14–19, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Gourdon:1993:COM

[GS93] Xavier Gourdon and Bruno Salvy. Computing one million digits of \sqrt{(}2). Maple Technical Newsletter, 0(10):66–71, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Guichard:1992:CNI

[Gui92] David R. Guichard. Counting non-isomorphic graphs with Maple. Maple Technical Newsletter, 0(8):52–56, Fall 1992. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn8.html.

Ceballos:1995:MRR

[GVC95] L. Gonzalez-Vega and L. Ceballos. Manipulation of real roots of polynomials: Isolating intervals or thom’s codes. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v2n2.html.

Hammer:1995:IDC

[Ham95] P. W. Hammer. Identification of diffusion coefficients in the heat equation using Maple. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n1.html.

Hausen:1994:GPL

[Hau94]  Jürgen Hausen.  Generat-

ing problems in linear algebra. Maple Technical Newsletter, 1(2):86–??, Fall 1994. ISSN 1061-5733. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtnv1n2.html.

Hausen:1995:GPL

[Hau95] J. Hausen. Generating problems in linear algebra II. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v2n2.html.

Holmes:1992:AMC

[HB92] Mark H. Holmes and Jonathan Bell. An application of Maple to chemical kinetics. Maple Technical Newsletter, 0(7):50–55, Spring 1992. ISSN 1061-5733. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn7.html.

Herod:1993:MHI

[Her93] James V. Herod. A model for an HIV infection. Maple Technical Newsletter, 0(10):72–78, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Herod:1997:TSD

[Her97] James V. Herod. Tutorial:

Solving differential equations using Maple. Maple Technical Newsletter, 4(2):72–77, 1997. ISSN 1061-5733.

Holland:1997:GII

[HK97] Shannon Holland and Matthias Kawski. 3D-graphics for iterated integrals. Maple Technical Newsletter, 4(1):92–98, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Harper:1994:DFE

[HM94] David Harper and C. D. Murray.

Disturbing function expansions.

Maple Technical Newsletter, 0: 24–28, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Herod:1997:NEV

[HM97] James Herod and Michael Monagan. A note from the Editors: Vol, 4, No. 2, 1997: Maple in the Physical Sciences. Maple Technical Newsletter, 4(2):1, 1997. ISSN 1061-5733.

Hoffmann:1994:MST

[Hof94] Christine W. U. Hoffmann. Modeling snow transport, an application of Maple. Maple Technical Newsletter, 1(1):69– 78, Spring 1994. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtnv1n1.html.

Heinle:1996:SAM

[HRR96] F. Heinle, R. Reng, and G. Runze. Symbolic analysis of multirate systems. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-

v3n1.html. Special Issue featuring Engineering Applications.

Howlett:1996:MVR

[HT96] C. Howlett and T. Tyhurst. The Maple V Release 4 graphical user interface. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/

maple/www/plibrary/mtn/mtn-
v3n2.html.

Herod:1994:MBM

[HY94] James V. Herod and Edward K.

Yeagers. Mathematics, biology, and Maple. Maple Technical Newsletter, 0:69–76, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Isbister:1994:QMM

[Isb94] Dennis Isbister. Quantum mechanics and Maple: The rotation matrix functions d_{mn}^{l}(β), 3j and 6-j symbols. Maple Technical Newsletter, 1(2):55–61, Fall 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n2.
html.

Jegou:1997:UMS

[Jég97] Sophie Jégou. Using Maple for symbolic differentiation to solve inverse problems. Maple Technical Newsletter, 4(2):32–40, 1997. ISSN 1061-5733.

Johnson:1992:WL

[Joh92]  Eugene Johnson.  Working

with linalg. Maple Technical Newsletter, 0(7):25–30, Spring 1992. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn7.html.

Karian:1994:RNG

[KG94] Zaven A. Karian and Rohit Goyal. Random number generation and testing. Maple Technical Newsletter, 1(1):32– 37, Spring 1994. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtnv1n1.html.

Krob:1997:ACM

[KLL^{+}97] D. Krob, A. Lascoux, B. Leclerc, J.-Y. Thibon, B. C. V. Ung, and S. Veigneau. Algebraic combinatorics with MAPLE and ACE. Maple Technical Newsletter, 4 (1):43–50, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Koepf:1996:SM

[Koe96]  W. Koepf.  Summation in

Maple. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v3n2.html.

Kraft:1994:MSR

[Kra94] Dieter Kraft. Modeling and simulating robots in Maple. Maple Technical Newsletter, 1(2):39– 47, Fall 1994. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtnv1n2.html.

Kloeden:1993:CSN

[KS93] Peter E. Kloeden and William D. Scott. Construction of stochastic numerical schemes through Maple. Maple Technical Newsletter, 0(10):60–65, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Kam:1998:IMP

[KTT98] Kiew M. Kam, Moses O. Tadé, and Benjamin L. C. To. Implementation of Maple procedures for simulating an industrial multi-stage evaporator. Maple Technical Newsletter, 5 (2–3):27–39, 1998. ISSN 1061-5733.

Labahn:1995:SLD

[Lab95]  G. Labahn.  Solving lin-

ear differential equations in Maple. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n1.html.

Lee:1996:MVR

[Lee96] T. Lee. Maple V Release 4: New features for engineers and scientists. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Lennartsson:1998:GCS

[Len98] Anders Lennartsson. On the generation of code for the solution of differential equations. Maple Technical Newsletter, 5 (2–3):95–101, 1998. ISSN 1061-

Lisser:1994:SKC

[Lis94] Bert Lisser. Solving Kostant’s conjecture using Maple. Maple Technical Newsletter, 0:29–34, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Layton:1996:IMM

[LK96] J. B. Layton and J. H. Kane. Integrating MathEdge with multimedia for instruction in engineering. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Lozano:1994:ITG

[LM94] Eugenio Roanes Lozano and Eugenio Roanes Macias. An implementation of “Turtle Graphics” in Maple V. Maple Technical Newsletter, 0:82–85, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Lopez:1994:SOE

[Lop94a] Robert J. Lopez. A saw-toothed oscillator: Exact and approximate solutions. Maple Technical Newsletter, 0:77–81, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Lopez:1994:SDE

[Lop94b] Robert J. Lopez. A separable differential equation: New insights. Maple Technical Newsletter, 1(1):79–82, Spring 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n1.
html.

Lopez:1996:TMI

[Lop96] R. Lopez. Tips for Maple instructors. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v3n2.html.

Low:1998:SFA

[Low98] K. H. Low. Symbolic frequency analysis of vibratory systems by the optimized Rayleigh method. Maple Technical Newsletter, 5 (2–3):67–75, 1998. ISSN 1061-5733.

Mansfield:1996:DAP

[Man96] E. Mansfield. The differential algebra package: diffgrob2. Maple Technical Newsletter, 3 (2):??, November 1996. ISSN 1061-5733. URL http://web.

mit.edu/afs/athena.mit.edu/
software/maple/www/plibrary/
mtn/mtn-v3n2.html.

Monagan:1992:OAD

[MD92] Michael M. Monagan and J. S. Devitt. The D operator and algorithmic differentiation. Maple Technical Newsletter, 0(7):17–24, Spring 1992. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn7.html.

Miller:1998:TPA

[MD98] John S. Miller and Michael J. Demetsky. Transportation planning applications: Using Maple for exploratory analysis. Maple Technical Newsletter, 5(2–3):76– 84, 1998. ISSN 1061-5733.

McPhedran:1992:BFI

[MDS92] R. C. McPhedran, D. H. Dawes, and Tony Scott. On a Bessel function integral. Maple Technical Newsletter, 0(8):33–38, Fall 1992. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn8.html.

Meade:1995:NGS

[Mea95]  D. B. Meade.  Numeri-

cal, graphical and symbolic analysis of Bernoulli equations. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v2n1.html.

Meade:1997:MPP

[Mea97] Douglas B. Meade. Mple and the Parachute Problem; modeling with an impact. Maple Technical Newsletter, 4(1):68–76, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Melancon:1997:FIW

[Mel97] Guy Melançon. Factorizing infinite words using Maple. Maple Technical Newsletter, 4(1):34– 42, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Meade:1996:STS

[MHW96] D. Meade, B. Haran, and R. White. The shooting technique for the solution of twopoint boundary value problems. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Mohseni:1995:EFM

[Moh95] P. Mohseni. EM field modeling using Maple. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n1.html.

Monagan:1988:AMC

[Mon88a] Michael Monagan. Announcement of Maple 4.0 for the Cray 2. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn2.html.

Monagan:1988:ASM

[Mon88b] Michael Monagan. Automatic simplification in Maple. Maple Newsletter, 0(3):??, July 1988. ISSN 1074-3790. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn3.html.

Monagan:1988:MPM

[Mon88c] Michael Monagan. On manipulating polynomials in Maple. Maple Newsletter, 0(3):??, July 1988. ISSN 1074-3790. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn3.html.

Monagan:1988:PFM

[Mon88d] Michael Monagan. A plotting facility in Maple. Maple Newsletter, 0(2):??, January 1988. ISSN 1074-3790. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn2.html.

Monagan:1989:CFM

[Mon89a] Michael B. Monagan. The collect function in Maple. Maple Newsletter, 0(4):??, January 1989. ISSN 1074-3790. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn4.html.

Monagan:1989:FTO

[Mon89b] Michael B. Monagan. Fortran translation and optimization. Maple Newsletter, 0(4):??, January 1989.

ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn4.html.

Monagan:1992:EEE

[Mon92] Michael B. Monagan. Exact eigenvalue and eigenvector computation in Maple. Maple Technical Newsletter, 0(8):24–32, Fall 1992. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn8.html.

Monagan:1993:TMU

[Mon93a] Michael B. Monagan. Tips for Maple users. Maple Technical Newsletter, 0(10):10–18, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Monagan:1993:UCA

[Mon93b] Michael B. Monagan. Using computer algebra to help understand the nature of eigenvalues and eigenvectors. Maple Technical Newsletter, 0(9):62– 68, Spring 1993. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn9.html.

Monagan:1994:MSA

[Mon94a] Michael Monagan. The Maple share (applications) library. Maple Technical Newsletter, 1 (1):14–16, Spring 1994. ISSN 1061-5733. URL http://www.

can.nl/Systems_and_Packages/

                                        Per_Purpose/General/Maple/
                                        mtn/mtnv1n1.html.

Monagan:1994:MVR

[Mon94b] Michael Monagan. Maple V Release 3. Maple Technical Newsletter, 1(1):7–13, Spring 1994. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtnv1n1.
                                        html.

Monagan:1994:TMUa

[Mon94c] Michael B. Monagan. Tips for Maple users. Maple Technical Newsletter, 1(1):17–23, Spring 1994. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtnv1n1.
                                        html.

Monagan:1994:TMUb

[Mon94d] Michael B. Monagan. Tips for Maple users. Maple Technical Newsletter, 1(2):11–18, Fall 1994. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtnv1n2.
                                        html.

Monagan:1995:MVR

[Mon95a] M. B. Monagan. The Maple V Release 4 math engine. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733.

                                        URL http://web.mit.edu/
                                        afs/athena.mit.edu/software/
                                        maple/www/plibrary/mtn/mtn-
                                        v2n2.html.

Monagan:1995:TMUa

[Mon95b] M. B. Monagan. Tips for Maple users and programmers. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n1.html.

Monagan:1995:TMUb

[Mon95c] M. B. Monagan. Tips for Maple users and programmers. Maple Technical Newsletter, 2(2):??, Fall 1995. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v2n2.html.

Monagan:1998:STI

[MS98] M. B. Monagan and M. A. Slawinski. The sensitivity of traveltime inversion for an anisotropic parameter in seismology. Maple Technical Newsletter, 5(2–3): 107–116, 1998. ISSN 1061-5733.

URL http://www.cecm.sfu.
ca/CAG/monagan_slawinski.
ps.gz.

Mulcahy:1996:BCS

[Mul96]  C. Mulcahy.  The basic

curves and surfaces of computer aided geometric design. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Murison:1997:ASO

[Mur97] Marc A. Murison. Analytical study of optical wavefront aberrations using Maple. Maple Technical Newsletter, 4(2):52– 62, 1997. ISSN 1061-5733.

McPhee:1996:ASA

[MW96] J. McPhee and C. Wells. Automated symbolic analysis of mechanical system dynamics. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Nagornov:1998:PAP

[NQC^{+}98] Oleg V. Nagornov, Nancy E. Quaranta, Marta G. Caligaris, Georgina B. Rodriguez, and Robert E. Caligaris. Propagation of air pollution in urban areas. Maple Technical Newsletter, 5(2–3):102–106, 1998. ISSN 1061-5733.

Ogunye:1998:NE

[OC98] Ayowale B. Ogunye and Tom Casselman. A note from the Editors. Maple Technical Newsletter, 5(2–3):1–2, 1998. ISSN 1061-5733.

Ogunye:1996:PCS

[Ogu96]  A. Ogunye.  Process con-

trol and symbolic computation: An overview with Maple V. Maple Technical Newsletter, 3(1):??, ???? 1996.

ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Ogunye:1998:SDT

[Ogu98]  Ayowale Ogunye.  Solution

of discrete-time symmetric matrix Diophantine equations using Maple V. Maple Technical Newsletter, 5(2–3):85–94, 1998. ISSN 1061-5733.

Ogilvie:1997:SCM

[OM97] J. F. Ogilvie and M. B. Monagan. Symbolic computation in molecular spectroscopy. Maple Technical Newsletter, 4(2):100– 108, 1997. ISSN 1061-5733.

Partin:1998:CAA

[Par98]  Lee R. Partin.  Combina-

torial analysis application for flowsheet synthesis of chemical plants. Maple Technical Newsletter, 5(2–3):15–26, 1998. ISSN 1061-5733.

Pathria:1991:ISM

[Pat91a] Anu Pathria. Indefinite summation in Maple. Maple Technical Newsletter, 0(5):??, Winter 1991. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn5.html.

Pathria:1991:ATB

[Pat91b] Anu K. Pathria. Archimedes’ theorem of the broken chord. Maple Technical Newsletter, 0 (6):??, Fall 1991. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn6.html.

Pennestri:1996:DCM

[PF96] E. Pennestri and V. Falasca. Design of cam mechanisms using Maple. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Piedboeuf:1996:MFR

[Pie96] J.-C. Piedboeuf. Modelling flexible robots with Maple. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-

v3n1.html. Special Issue featuring Engineering Applications.

Pintur:1994:MAD

[Pin94] David A. Pintur. MathEdge: The application development toolkit for Maple. Maple Technical Newsletter, 1(2):31–38, Fall 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n2.
html.

Pinchback:1996:MWW

[Pin96] R. M. Pinchback. Maple and the World Wide Web. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-

5733. URL http://web.mit.
edu/afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v3n2.html.

Pirastu:1995:AIS

[Pir95] R. Pirastu. Algorithms for indefinite summation of rational functions in Maple. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v2n1.html.

Porciello:1993:AMO

[PL93] G. Pino Porciello and Tom Lee.

Application of Maple in optimal control systems design and analysis. Maple Technical Newsletter, 0(10):79–84, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Promislow:1997:ACM

[Pro97] K. Promislow. Applications of center manifolds. Maple Technical Newsletter, 4(1):77–87, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Poblete:1997:ALL

[PVM97] Patricio V. Poblete, Alfredo Viola, and J. Ian Munro. Analyzing the LCFS linear probing hashing algorithm with the help of maple. Maple Technical Newsletter, 4(1):8–13, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Peikert:1991:PCS

[PWMdG91] R. Peikert, D. Wurtz, M. Monagan, and C. de Groot. Packing circles in a square. Maple Technical Newsletter, 0(6):??, Fall 1991. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn6.html.

RuizS:1996:UGB

[RF96] O. E. Ruiz S. and P. M. Ferreira. Using Gröbner bases in kinematic analysis of mechanisms. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Riel:1996:SSC

[Rie96] J. Riel. Syrup — A symbolic circuit analyzer. Maple Technical Newsletter, 3(1):??, ???? 1996. ISSN 1061-5733. URL http:

                                        //web.mit.edu/afs/athena.
                                        mit.edu/software/maple/www/
                                        plibrary/mtn/mtn-v3n1.html.

Special Issue featuring Engineering Applications.

Rybowicz:1996:TME

[RM96] M. Rybowicz and J.-P. Massia. Teaching Maple to engineering students. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-
v3n2.html.

Rough:1993:HAP

[RR93] A. D. Rough and J. Richardson. Harmonic analysis of phase controlled waveforms. Maple Technical Newsletter, 0(9):47– 53, Spring 1993. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn9.html.

Radia:1991:QAM

[RS91] Sanjay Radia and Tony C.

Scott. Queuing analysis of multiprocessing. Maple Technical Newsletter, 0(6):??, Fall 1991. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn6.html.

Riedel:1992:SVR

[RS92] Thomas Riedel and Prasanna Sahoo. Surfaces and volumes of revolution with Maple. Maple Technical Newsletter, 0(8):45– 51, Fall 1992. ISSN 1061-

5733. URL http://www.can.
nl/Systems_and_Packages/Per_
Purpose/General/Maple/mtn/
mtn8.html.

Richardson:1998:CPD

[RSPL98] Christina Richardson, Susan Sopher, Robert Petak, and Andreas Linninger. Conceptual process design using symbolic and numerical computing techniques. Maple Technical Newsletter, 5(2–3):3–14, 1998. ISSN 1061-5733.

Russell:1994:SWC

[Rus94] Geoff Russell. Shotgun wounding characteristics. Maple Technical Newsletter, 0:17–23, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Rybowicz:1992:ECC

[Ryb92] Marc Rybowicz. Extended capabilities for computing with algebraic functions and numbers. Maple Technical Newsletter, 0(8):16–23, Fall 1992. ISSN 1061-5733. URL http:

                                        //www.can.nl/Systems_and_
                                        Packages/Per_Purpose/General/
                                        Maple/mtn/mtn8.html.

Sacchetti:1997:BFL

[Sac97] Andrea Sacchetti. Band functions for the Lamé equation. Maple Technical Newsletter, 4 (1):28–33, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Salvy:1991:TWG

[Sal91]  Bruno Salvy.  Three ways

to get Stirling’s formula in Maple. Maple Technical Newsletter, 0(5):??, Winter 1991. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn5.html.

Scott:1993:RSB

[SBG93] T. C. Scott, Yehuda B. Band, and K. O. Geddes. Recipes for solving broad classes of definite integrals and applications. Maple Technical Newsletter, 0(10):19–27, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Schwalbe:1992:UPP

[Sch92] Daniel Schwalbe. Using phase portraits to understand population dynamics. Maple Technical Newsletter, 0(7):39–43, Spring 1992. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn7.html.

Scott:1988:NAD

[Sco88]  Tony Scott.  A new ap-

proach to definite integration for Maple 4.3. Maple Newsletter, 0(3):??, July 1988. ISSN 1074-3790. URL http:

//www.can.nl/Systems_and_
Packages/Per_Purpose/General/
Maple/mtn/mtn3.html.

Scott:1991:AVP

[Sco91a]  Tony Scott.  Application

of variational principles to a simple quantum mechanical problem using the Maple system. Maple Technical Newsletter, 0(5):??, Winter 1991. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn5.html.

Scott:1991:RSD

[Sco91b] Tony Scott. Recipes for solving definite integrals involving exponentials and logarithms (part II). Maple Technical Newsletter, 0(5):??, Winter 1991. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtn5.html.

Scott:1997:GAa

[Sco97a] Tony Scott. Guidelines for authors. Maple Technical Newsletter, 4(1):111, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Scott:1997:GAb

[Sco97b] Tony Scott. Guidelines for authors. Maple Technical Newsletter, 4(2):109, 1997. ISSN 1061-5733.

Scott:1994:AMM

[SFCM94] Tony Scott, Greg Fee, Rob Corless, and Mike Monagan. Applications of Maple to mathematical, scientific, and engineering problems. Maple Technical Newsletter, 0:49–57, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Scott:1997:GOF

[SGMS97] T. C. Scott, I. P. Grant, M. B. Monagan, and V. R. Saunders. Generation of optimized FOR-TRAN code for molecular integrals of Gaussian-type functions. Maple Technical Newsletter, 4(2):15–24, 1997. ISSN 1061-5733.

Stayman:1994:MIN

[SH94] J. W. Stayman and Mark H.

Holmes. Maple: as an interface for numerically solving chemical kinetic equations. Maple Technical Newsletter, 1(2):48–54, Fall 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n2.
html.

Shahidi:1997:SNT

[Sha97] Reza Shahidi. A short note on trees, foxes and Maple. Maple Technical Newsletter, 4(2):86– 92, 1997. ISSN 1061-5733.

Taylor:1996:SSD

[SKT96] D. Schwalbe, H. Kooijman, and R. Taylor. Solving stiff differential equations and differential algebraic systems with Maple V. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v3n2.html.

Scott:1994:NEa

[SM94] Tony Scott and Michael Monagan. A note from the Editors. Maple Technical Newsletter, 0:1, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Scott:1994:MSE

[SMC94] Tony Scott, Blair Madore, and Robert M. Corless. Maple in science education. Maple Technical Newsletter, 0:58–68, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Scott:1991:AMM

[SMFC91] Tony C. Scott, Michael B. Monagan, Greg Fee, and Robert Corless. Applications of Maple to mathematical, scientific and engineering problems. Maple Technical Newsletter, 0(6):??, Fall 1991. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn6.html.

Scott:1994:MEP

[SMPR94] Tony Scott, Michael B. Monagan, Richard Pavelle, and Darren Redfern. Maple in education — Part II. Maple Technical Newsletter, 1(1):69–78, Spring 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n1.

html. See also [SPR92].

Scott:1993:NEb

[SP93] Tony Scott and Reid M. Pinchback. A note from the Editors. Maple Technical Newsletter, 0(10):1–5, Fall 1993. ISSN 1061-5733.

Scott:1994:NEb

[SP94] Tony Scott and Reid M. Pinchback. A note from the Editors. Maple Technical Newsletter, 1(1):1–3, Spring 1994. ISSN 1061-5733.

Pavele:1992:ME

[SPR92] Tony Scott, Richard Pavele, and Darren Redfern. Maple in education. Maple Technical Newsletter, 0(7):11–16, Spring 1992. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn7.html.

See also [SMPR94].

Scott:1992:NEa

[SR92a] Tony Scott and Darren Redfern.

A note from the Editors. Maple Technical Newsletter, 0(7):1–3, Spring 1992. ISSN 1061-5733.

Scott:1992:NEb

[SR92b] Tony Scott and Darren Redfern.

A note from the Editors. Maple Technical Newsletter, 0(8):1–3, Fall 1992. ISSN 1061-5733.

Scott:1993:NEa

[SR93] Tony Scott and Darren Redfern.

A note from the Editors. Maple Technical Newsletter, 0(9):1–3, Spring 1993. ISSN 1061-5733.

Halling:1995:CPD

[SRGH95] Z. Shung-Ren, J. Grotendorst, and H. Halling. Calculation of the potential distribution for a three-layer spherical volume conductor. Maple Technical Newsletter, 2(1):??, Spring 1995. ISSN 1061-5733. URL http:

//web.mit.edu/afs/athena.
mit.edu/software/maple/www/
plibrary/mtn/mtn-v2n1.html.

Scott:1997:AMN

[STB97] T. C. Scott, M. Trippenbach, and Y. B. Band. An application of Maple to near-field optics. Maple Technical Newsletter, 4(2):63–71, 1997. ISSN 1061-5733.

Stockie:1998:GIT

[Sto98] John M. Stockie. The geometry of intersecting tubes applied to controlling a robotic welding torch. Maple Technical Newsletter, 5(2–3):58–66, 1998. ISSN 1061-5733.

Strehl:1993:BSI

[Str93] Volker Strehl. Binomial sums and identities. Maple Technical Newsletter, 0(10):37–49, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

Taylor:1994:TMV

[Tay94]  Ross Taylor.  Thermody-

namics with Maple V II, phase equilibria in binary systems. Maple Technical Newsletter, 1(1):83–92, Spring 1994. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtnv1n1.

html. See also [TM93, AT94, Tay96].

Taylor:1996:TMI

[Tay96] R. Taylor. Thermodynamics with Maple: IV, the properties of steam. Maple Technical Newsletter, 3(2):??, November 1996. ISSN 1061-5733.

URL http://web.mit.edu/
afs/athena.mit.edu/software/
maple/www/plibrary/mtn/mtn-

v3n2.html. See also [TM93, Tay94, AT94].

Taylor:1993:TME

[TM93] Ross Taylor and Michael B.

Monagan. Thermodynamics with Maple: I, equations of state. Maple Technical Newsletter, 0(10):50–59, Fall 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn10.html.

See also [Tay94, AT94, Tay96].

Vahle:1993:SC

[Vah93]  Michael O. Vahle.  Solv-

ing the congruence x^{2} ≡ a mod n. Maple Technical Newsletter, 0(9):69–76, Spring 1993. ISSN 1061-5733. URL

http://www.can.nl/Systems_
and_Packages/Per_Purpose/
General/Maple/mtn/mtn9.html.

Wald:1997:SUP

[Wal97] Burkhard Wald. The semigroup of unary pairfunctions, verification of the group relations. Maple Technical Newsletter, 4 (1):51–54, Winter 1997. ISSN 1061-5733. Special issue on Maple in the mathematical sciences.

Wurtz:1994:HPC

[WMP94] D. Würtz, M. Monagan, and R. Peikert. The history of packing circles in a square. Maple Technical Newsletter, 0:35–42, 1994. ISSN 1061-5733. Maple in Mathematics and the Sciences — A Special issue.

Zimmermann:1994:GPR

[Zim94]  Paul Zimmermann.  Gaïa:

A package for the random generation of combinatorial structures. Maple Technical Newsletter, 1(1):38–46, Spring 1994. ISSN 1061-5733. URL

                                        http://www.can.nl/Systems_
                                        and_Packages/Per_Purpose/
                                        General/Maple/mtn/mtnv1n1.
                                        html.